# EngChain: A Symbolic Benchmark for Verifiable Multi-Step Reasoning in Engineering

### v1

@@ Appendix I LLM-as-Judge Error Categories
Table 7: Overview of all models evaluated, grouped by functionality: (1) frontier proprietary LLMs representing current performance upper bounds, (2) science/engineering-enhanced models trained on technical corpora, (3) math-enhanced models for quantitative reasoning, and (4) general-purpose open LLMs providing broad foundations. For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
Appendix L AI-Assisted QA Workflow: Template Examples
This appendix provides visual evidence of the LLM Peer Reviewer’s output for an automatically approved template (provided in Figure 9) and a flagged template (provided in Figure 10), illustrating the function of our automated Quality Assurance (QA) process (Section 4.5).
Figure 9: LLM Peer Review result for an Approved Template (template_rackett_equation_volume). The template, designed for Chemical Engineering Thermodynamics, receives the maximum score of 5/5 across all core categories (Physical Plausibility, Mathematical Correctness, and Pedagogical Clarity), indicating that the problem structure is robust and ready for benchmark inclusion.
Figure 10: LLM Peer Review result for a Flagged Template (template_euclidean_distance_binary). The model detected a fundamental flaw, assigning a score of 3/5 for both Plausibility and Correctness. The explicit explanation flags a “calculation error that sets sqrt(Eb) to zero,” which immediately forces the template out of the automated pipeline and into human review for manual correction.
Appendix M Detailed LLM-as-Judge Error Analysis
This appendix provides the full, fine-grained breakdown of the qualitative error analysis performed on Gemini 2.5 Pro, using GPT-4o as the judge. The following tables detail the distribution of all 4,118 reasoning steps that were flagged by the quantitative verifier as a mismatch from the ground-truth solution. Table 8 breaks down the errors by engineering branch,  Table 9 shows the distribution by problem difficulty, and Figure 11 provides a proportional breakdown of error types across all nine engineering domains.
Branch Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Chemical 1353 206 67 17 43 1686
Electrical 827 221 23 24 94 1189
Mechanical 865 207 63 40 67 1242
Total 3045 634 153 81 204 4117
Table 8: Branch-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three main engineering branches.
Difficulty Alternative Correct Conceptual Error Calculation Error Input Error Other Total
Easy 1088 162 52 15 61 1378
Intermediate 1178 204 65 51 73 1571
Advanced 779 268 36 15 70 1168
Total 3045 634 153 81 204 4117
Table 9: Difficulty-level Error Analysis for Gemini 2.5 Pro. This table shows the distribution of 4,117 flagged reasoning mismatches, categorized by the judge (GPT-4o), across the three problem difficulty levels.
Figure 11: Domain-level Error Analysis for Gemini 2.5 Pro. This 100% stacked bar chart illustrates the proportional distribution of error types across all nine engineering domains. This visualization highlights how the ratio of error types changes by domain; for instance, Fluid Mechanics and Signals & Systems show a much larger proportion of Conceptual Errors and Other errors compared to Thermodynamics or Reaction Kinetics, where Alternative Correct solutions dominate.

### v2

@@ Appendix I Human Certification Interface
Table 12: Overview of the 24 models evaluated, grouped into (1) frontier proprietary, (2) general-purpose open, and (3) math-enhanced categories. For proprietary models, we list the internal or API model identifiers; for open models, we include the HuggingFace source.
Appendix O Inference Prompt and Output Processing
O.1 Standardized Inference Prompt
To ensure consistent and parsable outputs across all models, we employed a standardized zero-shot system prompt. This prompt explicitly instructs the model to structure its response with specific headings and to format the final result for automated extraction.
You are an expert engineer. Solve the following
problem by providing a detailed, structured
solution. Use the exact headings and
formatting provided below.
# Given
List all known variables and their
values with units.
# Find
State the variable(s) to be calculated.
# Formulae
Write down all necessary governing equations
before substituting any values.
# Solution
Provide a step-by-step calculation. Each step
must start on a new line and be formatted exactly
as ’**Step X:**’, where X is the step number.
Show the substitution of values into
the formulae clearly.
# Final Answer
State the final numerical result with its units
in the format: **Answer:** [value] [units]
O.2 Output Parsing Logic
To process the model outputs generated by the prompt above, we employ a heuristic parsing pipeline:
  1. 1.
    Step Segmentation: We use regular expressions to identify line-start markers matching either the Step <number> format (e.g., **Step 1**) or the numbered-list format (e.g., 1.). A subsequent regex cleans these prefixes to isolate the reasoning text.
  2. 2.
    Value Extraction: We extract numerical values from each step and the final answer using a three-priority heuristic:
    • Priority 1: Matching the explicit **Answer:** tag defined in the prompt.
    • Priority 2: Identifying the last number following an equals sign (e.g., V = 100.0).
    • Priority 3: Extracting the last standalone number in the segment.
  3. 3.
    Fallback: If no overall **Answer:** tag is found, the final numerical value extracted from the last reasoning step is used for verification.